Fermi National Laboratory
 

Operation Summary for May 27, 2011 - May 31, 2011

The Integrated Luminosity for the period from 5/23/11 to 5/30/11 was 58.41 pb-1 (inverse picobarns).

Friday May 27

The day shift began with Operations monitoring the Tevatron (TEV) store 8768 (154.83E30) and the Recycler stash (182.7E10), with the Antiproton Source (Pbar) stacking (23.6E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off due to target replacement, with the MiniBooNE and Meson FTBF T978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8768 at 1:30 PM. Operations reported that the Tevatron Electron Lens (TEL1) current became noisy during the store, which caused minor beam problems. Cryo system personnel began the TEV sector B3 wet engine repairs. TEV personnel started 150 GeV studies during the repairs. Operators reported that the C1 BLM VME crate had failed and that controls personnel replaced the crate. Cryo system personnel tested the TEV A0 nitrogen valves.

During the evening shift, Operations disabled beam to T-978 at 5:51 PM. Cryo system personnel continued with the TEV B3 wet engine repairs; later they reported that the wet engine required a complete overhaul, but the repairs would take quite a few hours. Since the B3 cryo system had spare cryo capacity, the Cryo Coordinator decided to try to ramp the TEV to 980 GeV using the JT valve instead of the wet engine. The B3 cryo system survived the ramp to 980 GeV. The Run Coordinator agreed to try to establish a store in this mode and operators put the TEV into HEP shot setup. Operations established store 8770 at 6:59 PM with an Average Initial Luminosity (AIL) of 366.69E30. Operators reported that a Pbar Accumulator vertical correction magnet power supply had failed; operators replaced the supply. Operators tuned the Booster to correct beam positions.

Saturday May 28

The midnight shift began with Operations monitoring store 8770 (154.81E30) and the stash (146.6E10), with Pbar stacking (12.1E10), and with MiniBooNE taking beam.

During the midnight shift, operators monitored the B3 cryo system all shift. Operations established beam to T-978 at 4:48 AM.

During the day shift, at 8:26 AM, store 8770 quenched. Operators traced the cause to a pre-fire of a TEV antiproton abort kicker (C:AAK5). TEV kicker personnel investigated and lowered the reservoir voltage. Operators reported that the TEV A1 cryo system cooldown stalled. Cryo system personnel took the rest of the shift to cool down the system. Operations prepared the CDF and D0 collision halls for access. Operations disabled beam to T-978 at 6 PM.

During the evening shift, Operators reported at 8:53 PM that the TEV A1 cryo system was cold. Operators conducted a TEV turn on, orbit check, and put the TEV into shot setup at 10:32 PM. Operations established store 8772 at 11:36 PM with an AIL of 403.25E30.

Sunday May 29

The midnight shift began with Operations monitoring store 8772 and the stash, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, the MI-52 kicker tripped off; operators had to reset the power supply locally. Operations established beam to T-978 at 6:32 AM. At 7:28 AM, the MI power supply interlock loop tripped off on a ground fault indication during a transfer of antiprotons; the transfer was lost (about 12E10).

During the day shift, Operations terminated store 8772 at 8:07 AM. Operators prepared the D0 collision hall for a short access. Operations put the TEV into shot setup at 10:20 AM, however a lightning strike caused the TEV ramp to dump. After operators put the TEV through a turn around and check out, they put the TEV into shot setup at 1:26 PM. Operations established store 8777 at 1:23 PM with an AIL of 373.12E30. Lightning strikes caused several short down times due to the Booster and Linac RF systems.

During the evening shift, Operations disabled beam to T-978 at 6 PM. Operators reported other short down times due to problems with the MI-52 kicker and a Linac quadrupole power supply (QPS130). Operators accessed the H- Source to investigate the loss of source pressure. Operators increased the hydrogen bottle pressure and recovered the source. This caused no problems since the I- Source was supplying beam to the Linac.

Monday May 30

The midnight shift began with Operations monitoring store 8777, with Pbar stacking, and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8777 at 12:35 AM. Operators prepared the TEV for a quick access to investigate a high water sump alarm at C3. Operators and FESS personnel accessed and found that the pump was working correctly, but was having trouble keeping up due to the large amount of rain that fell earlier in the day. Operations established beam to T-978 at 4:30 AM. Operations put the TEV into shot setup at 5:54 AM and established store 8781 at 7:13 AM with an AIL of 378.21E30. Operators reported minor down times caused by a vacuum burst in the MI-8 line and another trip of the MI-52 kicker.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI personnel to conduct a study. MTA might be taking beam later in the day.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/31/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory